# 1D Spontaneous Symmetry Breaking in thermal equilibrium via Non-Hermitian Construction

### v1

@@ Abstract
Introduction— Spontaneous symmetry breaking (SSB) is a central concept in particle and condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8], where the low-energy physics displays lower symmetry than the parent Hamiltonian, e.g., a ferromagnet, despite the parent Hamiltonian being spin symmetric, or a superconductor with broken ‘gauge’ symmetry, despite the parent Hamiltonian being manifestly gauge invariant. The Mermin-Wigner theorem has pointed out that due to proliferating defects and domain walls, SSB and long-range order do not materialize in one-dimensional (1D) systems with local interactions in thermal equilibrium [1, 2, 3, 8, 9, 10, 11, 12]. On the other hand, exceptions exist in classical systems far from thermal equilibrium, as SSB in 1D dynamical, diffusive and driven systems [13, 14, 15, 16, 17].
@@
There has been much recent progress on non-Hermitian quantum systems, such as non-Hermitian topological phenomena [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], the non-Hermitian skin effect (NHSE) [30, 32, 33, 34, 31], and nontrivial path-integral winding [35, 36]. While quantum mechanics dictates Hermitian Hamiltonians, the non-Hermitian models are increasingly proving their relevance in the fields of condensed matter physics [37, 38], finding their origins in open or dissipative systems [39, 40, 41, 42, 43, 44, 45, 46], optical [47, 48, 49, 50, 51, 52], acoustic [53, 54, 55, 56, 57, 58, 59, 60], electric [61, 62, 63, 64, 65, 66], mechanical [67, 68, 69, 70, 71, 72], and solid-state systems [73]. Recently, non-Hermitian physics of non-reciprocal classical fields has laid the foundations for understanding SSB in dynamical and non-equilibrium systems [17].
@@
Here, we re-visit SSB in 1D quantum systems from the perspective of non-Hermitian constructions. Without loss of generality, we study 1D quantum systems where non-Hermitian fermions are locally coupled to a discrete field. The overall quantum systems are fully Hermitian and efficient for numerical and sign-problem-free Monte Carlo (MC) calculations. Interestingly, despite their 1D nature and local interactions, such systems exhibit clear signatures of long-range order following SSB, such as order parameters and specific heat behaviors, as the temperature is lowered. Therefore, for the first time, we deliver a 1D SSB scenario, which, unlike previous examples, is under thermal equilibrium and beyond classical dynamics.
@@
The reasons that our constructions enable 1D SSB with local interactions and thermal equilibrium are as follows. From a energetic perspective, non-Hermitian systems exhibit unique non-locality - the NHSE shows that boundaries or domain walls impact their physics globally [30, 32, 33, 34, 31], and non-Hermitian delocalization yields long-range correlations and entanglements [35, 36] - therefore, non-Hermitian constructions may effectively introduce attractive strings between domain walls that contain their proliferation. From a fermionic perspective, the phase transition accompanies a characteristic and essentially topological shift of fermion path integrals from conventional ($w=0$) to nontrivial (simultaneous $w=\pm 2,\pm 4,\cdots$) path-integral winding-number sectors across the periodic boundary conditions (PBCs). Results from numerical calculations and mean-field theories also portray consistent physics. Although overall Hermitian, various non-Hermitian properties, e.g., such as nonzero winding number and nontrivial Greens functions [74], emerge in the ordered phase, exhibiting what we dub spontaneous ‘Hermiticity’ breaking - a novel route toward non-Hermitian physics.
Models— Without loss of generality, we consider the following 1D Hamiltonian under PBCs:
@@
where $c_{i}$ is a fermion annihilation operator at site $x$, and $X_{i}=\pm 1$ takes discrete values on the bond between sites $i$ and $i+1$. $t=1$ and $t^{\prime}$ are the nearest-neighbor and next-nearest-neighbor hopping amplitudes of the fermions, respectively, $J$ is an Ising interaction between the nearest-neighbor $X_{i}$, and $U$ describes a local coupling between the fermions and $X_{i}$; see Fig. 1 for illustration.
We consider the property of $\hat{H}$ in thermal equilibrium at temperature $\beta=1/k_{B}T$, represented by its density operator $\hat{\rho}\propto e^{-\beta\hat{H}}$ and partition function:
$Z=\sum_{X}e^{-\beta E_{J}(X)}\cdot\mbox{Tr}_{c}[e^{-\beta\hat{H}(X)}]=\sum_{X}e^{-\beta E_{J}(X)}\prod_{n}[{1+e^{-\beta\varepsilon_{n}(X)}}],$ (2)
where $E_{J}(X)=-J\sum_{i}X_{i}X_{i+1}$, and $\varepsilon_{n}(X)$ is the spectrum of non-Hermitian fermion model $\hat{H}(X)$ following a specific ${X}$ configuration. For physical quantities’ expectation values, we average over $X$ with the probability:
$p({X})=\frac{1}{Z}e^{-\beta E_{J}(X)}\prod_{n}[{1+e^{-\beta\varepsilon_{n}(X)}}],$ (3)
which, given $\varepsilon_{n}(X)$ appears in complex-conjugate pairs, remains positive-definite and allows sign problem-free MC sampling. Further, through a similarity transformation, $\hat{H}(X)$ for $t^{\prime}=0$ maps to a translation-invariant Hatano-Nelson (HN) model, which depends only on the number $n_{\pm}$ of $X_{i}=\pm 1$ bonds; consequently, the fermion spectrum and weights are highly degenerate and the summation in Eq. 2 can be carried out numerically and exactly. Alternatively, we can analyze the model qualitatively via the self-consistently mean-field theory [74]:
$\hat{H}_{c}$ $=$ $\sum_{k}\left[2t\cos(k)+2t^{\prime}\cos(2k)+2iU\langle X\rangle_{\beta}\sin(k)\right]c^{\dagger}_{k}c_{k},$
$\hat{H}_{X}$ $=$ $X\left[\frac{2iU}{L}\sum_{k}\langle c^{\dagger}_{k}c_{k}\rangle_{\beta}\sin(k)-2J\langle X\rangle_{\beta}\right],$ (4)
where we have neglected the fluctuations of $X$ away from $\langle X\rangle_{\beta}$, and the $\langle\cdots\rangle_{\beta}$ are thermal expectation values. For further details of methods, see the Supplemental Material [74].
@@
Despite non-Hermitian $\hat{H}(X)$, the overall system remains Hermitian as we sum over $X$ configurations:
where we have used $\hat{H}(-X)=\hat{H}^{\dagger}(X)$ and thus $\hat{\rho}(-X)=\hat{\rho}^{\dagger}(X)$ for the fermions given the $X$ configuration. The model is also invariant under translation as well as $X\rightarrow-X$ and $i\rightarrow-i$.
Figure 2: The order parameter $\langle X\rangle$ and specific heat $C_{V}$ indicate the presence of SSB and long-range order in the 1D models in Eq. 1 below a finite critical temperature $\beta>\beta_{C}$. The vertical dashed lines labels $\beta_{C}$ we pinpoint for $t^{\prime}=0.0$ through finite-size analysis (inset). Here, $U=0.4$, $J=0.0$, $t^{\prime}=0.0,0.5$, and $L=70$. We provide consistent MC and exact results for $t^{\prime}=0$, whereas only MC results are available at $t^{\prime}=0.5$.
Long-range order and energetics— We numerically evaluate physical properties such as the expectation value $\langle X\rangle$ and the specific heat $C_{V}$ over various system sizes and observe clear signatures of ordered phases at sufficiently low temperatures ($\beta>\beta_{C}$); see examples in Fig. 2. Here, $\langle X\rangle$ serves as an Ising-type order parameter and its long-range correlation emerges in the ordered phases [74]. The finite-size scaling of the specific heat $C_{V}$ accurately locates the transitions, which are essentially first order, as the preferential $\langle X\rangle$ the minimum of free energy $F(\langle X\rangle)$ jumps from one value to another without continuity [74]. We summarize the results in the phase diagrams in Fig. 3, where the ordered phases hold finite regions leaning towards larger $U$ and lower $T$, indicating their sufficient robustness towards variations of parameters. We note that aside from the first and ordering transition, there exists a series of additional transitions, whose physics we will explain later. On the contrary, the entire parameter spaces consist of only a disordered phase under open boundary conditions (OBCs) or without non-Hermitian construction, e.g., $U=0$ or $U=0.4i$ [74].
@@
Let us consider the $t^{\prime}=0$ case in Eq. 1 as an example: given a specific $X$ configuration, the ground-state energy $E(X)$ of $\hat{H}(X)$ depends only on $n_{\pm}$; the minimum $E(X)$ corresponds to the uniform configuration $X_{0}$ with all $X_{i}=1$ ($X_{i}=-1$). Importantly, unlike the Ising interaction $E_{J}(X)$ that contributes a constant energy cost of $2J$ for each domain wall, $E(X)$, through non-Hermitian construction, is no longer locally dependent on the (number of) domain walls in $X$, but on the separations between them as well. For instance, $E(X)$ increases linearly with two domain walls $r=\alpha L$ apart, while $E(X)$ saturates to a constant with nearby domain walls in the thermodynamic limit, as we illustrate in Fig. 4. The trend remains at finite temperature. Such non-local energetic dependence applies to general scenarios beyond similarity transformations, e.g., $t^{\prime}\neq 0$, but are absent in similar yet Hermitian constructions or under OBCs [74].
Figure 4: The ground-state energy $E(X)$ increases linearly with two domain walls $r=\alpha L$ apart and saturates to a constant with nearby domain walls in the thermodynamic limit. Note the differences between the vertical scales. Here, $t^{\prime}=0,0.5$, $U=0.4$, and $X_{0}$ is uniformly ordered without domain walls. The insets illustrate the schematics of the corresponding $X$ configurations with two domain walls strung together by $E(X)$ and the constant energy cost $2J$ from $E_{J}$.
@@
Topological aspects of fermion path integrals— We look into the fermion path integral in 1+1-dimensional space-(imaginary-)time underlying such SSB and ordered phases, once the $X$ field in Eq. 1 is integrated out. The consequences look trivial at first glimpse, as the periodicity in imaginary time $\beta$ mandates the fermions always to traverse an equal number of strengthened and attenuated bonds, and numerical results indeed eliminate the existence of nontrivial physics such as SSB or order phase under OBCs despite identical non-Hermitian construction [74]. Importantly, however, the fermion trajectories may circle the systems under PBCs, a topological nature labeled by an integer path-integral winding number $w$. The possibility of nonzero $w$ has been long overlooked, as the most relevant path integrals predominantly concentrate in the $w=0$ sector in conventional Hermitian models [77]. More recently, $w\neq 0$ sectors have begun to play vital roles, but only in non-Hermitian models. Nevertheless, as Hermitian conjugate maps $w$ to $-w$, having path integral simultaneously concentrated in the ±$\pm w$ sectors is equally valid in Hermitian quantum systems. Indeed, our phase diagram is characterized by such a topological nature: the SSB is signaled by the transition from the $w=0$ sector to the $w\neq 0$ sectors.
@@
Figure 5: The phases and phase transitions exhibit a topological nature from the fermion perspective: its fermion path integral’s winding number $w=\operatorname{Im}(\langle v\rangle)\beta/L$ evolve from the conventional $w=0$ plateau in the disordered phase to nontrivial plateaus of $w=\pm 2,\pm 4,\cdots$ in the ordered phase. Here, $U=0.4$, $J=0.0$, and $t^{\prime}=0.0$. The vertical dashed line labels the corresponding transitions in Fig. 2. Inset: illustrations of trivial $w=0$ and nontrivial $w=2$ winding numbers in the fermion path integral under PBCs. Lower left: the probability distribution of $v$ shifts from peaking at 0 to finite-imaginary-value concentrations. Note the symmetry upon $v\rightarrow-v$, as Hermiticity requires. Lower right: long-range correlation of $v$ emerges in the ordered phases. The respective temperatures in the lower panels correspond to the colored arrows in the upper panel.
[…]
The SSB also leaves nontrivial marks on the fermions’ spectral functions: while singularities mainly concentrate on the real energy axis in the disordered phase, they emerge in imaginary regions on the complex energy plane in the order phases [74]. Such complex spectral properties, as well as the finite $\operatorname{Im}(\langle v\rangle)\neq 0$ in one of the SSB sectors, commonly relate to non-Hermitian physics - namely, such ordered systems have a spontaneously broken “Hermiticity”  [^4] The concept of spontaneous ‘Hermiticity’ breaking may sound unphysical at first glimpse. However, an analogy can be drawn with the broken ‘gauge’ symmetry in a superconducting phase - both Hermiticity and gauge invariance are fundamental requirements of a quantum theory and broken only at a mean-field level within a particular order-parameter sector, equivalent to other sectors via transformations, i.e., gauge transformations or Hermitian conjugate. . Such non-Hermitian features are also evident from a mean-field perspective, e.g., Eq. 4 with a finite $\langle X\rangle_{\beta}$.
@@
Acknowledgment: We acknowledge insightful discussions with Yongxu Fu and generous support from the National Natural Science Foundation of China (Grants No.92270102 & No.12174008) and the National Key R&D Program of China (Grant No.2022YFA1403700).

### v2

@@ Abstract
Introduction— Spontaneous symmetry breaking (SSB) is a central concept in particle and condensed matter physics [1, 2, 3, 4, 5, 6, 7, 8], where the low-energy physics displays lower symmetry than the parent Hamiltonian, e.g., a ferromagnet, despite the parent Hamiltonian being spin symmetric, or a superconductor with broken ‘gauge’ symmetry, despite the parent Hamiltonian being manifestly gauge invariant. Due to proliferating defects and domain walls, SSB and long-range order do not materialize in one-dimensional (1D) systems with local interactions in thermal equilibrium [1, 2, 3, 8, 9, 10, 11, 12]. On the other hand, exceptions exist in classical systems far from thermal equilibrium, as SSB in 1D dynamical, diffusive and driven systems [13, 14, 15, 16, 17].
@@
There has been much recent progress on non-Hermitian quantum systems, such as non-Hermitian topological phenomena [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31], the non-Hermitian skin effect (NHSE) [30, 32, 33, 34, 31], and nontrivial path-integral winding [35, 36]. Another prominent topic is the $\mathcal{P}\mathcal{T}$-symmetry breaking in non-Hermitian physics, characterized by a spectral transition from purely real eigenvalues to complex-conjugate pairs at critical parameter values [37, 38, 39, 40]. While quantum mechanics dictates Hermitian Hamiltonians, the non-Hermitian models are increasingly proving their relevance in the fields of condensed matter physics [41, 42], finding their origins in open or dissipative systems [43, 44, 45, 46, 47, 48, 49, 50], optical [51, 52, 53, 54, 55, 56], acoustic [57, 58, 59, 60, 61, 62, 63, 64], electric [65, 66, 67, 68, 69, 70], mechanical [71, 72, 73, 74, 75, 76], and solid-state systems [77]. Recently, non-Hermitian physics of non-reciprocal classical fields has laid the foundations for understanding SSB in dynamical and non-equilibrium systems [17].
@@
Here, we re-visit SSB in 1D quantum systems from the perspective of non-Hermitian constructions. We study 1D quantum systems where non-Hermitian fermions are locally coupled to a discrete field. The overall quantum systems are fully Hermitian and efficient for numerical and sign-problem-free Monte Carlo (MC) calculations. Interestingly, despite their 1D nature and local interactions, such systems exhibit clear signatures of long-range order following SSB, such as order parameters and specific heat behaviors, as the temperature is lowered. Therefore, for the first time, we deliver a 1D SSB scenario, which, unlike previous examples, is under thermal equilibrium and beyond classical dynamics.
@@
The reasons that our constructions enable 1D SSB with local interactions and thermal equilibrium are as follows. From a energetic perspective, non-Hermitian systems exhibit unique non-locality - the NHSE shows that boundaries or domain walls impact their physics globally [30, 32, 33, 34, 31], and non-Hermitian delocalization yields long-range correlations and entanglements [35, 36] - therefore, non-Hermitian constructions may effectively introduce attractive strings between domain walls that contain their proliferation. From a fermionic perspective, the phase transition accompanies a characteristic and essentially topological shift of fermion path integrals from conventional ($|w|=0$) to nontrivial (simultaneous $|w|=2,4,\cdots$) path-integral winding-number across the periodic boundary conditions (PBCs). Results from numerical calculations and mean-field theories also portray consistent physics. Although overall Hermitian, various non-Hermitian properties, e.g., such as nonzero winding number and nontrivial fermion spectral functions [78], emerge in the ordered phase, exhibiting what we dub spontaneous ‘Hermiticity’ breaking - a novel route toward non-Hermitian physics.
Models— We consider the following 1D Hamiltonian under PBCs:
@@
where $c_{i}$ is a fermion annihilation operator at site $i$, and $X_{i}=\pm 1$ takes discrete values on the bond between sites $i$ and $i+1$. $t=1$ and $t^{\prime}$ are the nearest-neighbor and next-nearest-neighbor hopping amplitudes of the fermions, respectively, $J$ is an Ising interaction between the nearest-neighbor $X_{i}$, and $U$ describes a local coupling between the fermions and $X_{i}$; see Fig. 1 for illustration.
The thermal equilibrium properties of the system at temperature $\beta=1/k_{B}T$ are described by the density matrix $\hat{\rho}\propto e^{-\beta\hat{H}}$. For a specific $X$ configuration, the Hamiltonian $\hat{H}(X)$ is non-Hermitian, yet the overall system remains Hermitian as we sum over $X$ configurations:
where we have used $\hat{H}(-X)=\hat{H}^{\dagger}(X)$ and thus $\hat{\rho}(-X)=\hat{\rho}^{\dagger}(X)$ for the fermions given the $X$ configuration. This derivation implicitly relies on the system’s symmetry $D\rho D^{-1}=\rho$ under the configuration transformation $D:X\to-X$, which will spontaneously break at the phase transition.
@@
The partition function takes the form:
$Z=\sum_{X}\mbox{Tr}_{c}[e^{-\beta\hat{H}(X)}]=\sum_{X}\prod_{n}[{1+e^{-\beta\varepsilon_{n}(X)}}],$ (3)
where $\varepsilon_{n}(X)$ denotes the eigenvalues of $\hat{H}(X)$, which are either purely real or form complex-conjugate pairs due to the $\mathcal{P}\mathcal{T}$ symmetry of $\hat{H}(X)$[78]. For physical quantities’ expectation values, we average over $X$ with the probability:
$p({X})=\frac{1}{Z}\prod_{n}[{1+e^{-\beta\varepsilon_{n}(X)}}],$ (4)
which remains positive-definite (due to the complex-conjugate-pair spectrum) and circumvents the MC sign problem. Further, through a similarity transformation, $\hat{H}(X)$ for $t^{\prime}=0$ maps to a translation-invariant Hatano-Nelson (HN) model, which depends only on the number $n_{\pm}$ of $X_{i}=\pm 1$ bonds; consequently, the fermion spectrum and weights are highly degenerate and the summation in Eq. 3 can be carried out numerically and exactly [78]. Alternatively, we can analyze the model qualitatively via the self-consistently mean-field theory:
$\hat{H}_{c}$ $=$ $\sum_{k}\left[2t\cos(k)+2t^{\prime}\cos(2k)+2iU\langle X\rangle\sin(k)\right]c^{\dagger}_{k}c_{k},$
$\hat{H}_{X}$ $=$ $X\left[\frac{2iU}{L}\sum_{k}\langle c^{\dagger}_{k}c_{k}\rangle\sin(k)-2J\langle X\rangle\right],$ (5)
where we have neglected the fluctuations of $X$. For further details of methods, see the Supplemental Material [78].
Figure 2: The order parameter $\langle X\rangle$ and specific heat $C_{V}$ indicate the presence of SSB and long-range order in the 1D models in Eq. 1 below a finite critical temperature $\beta>\beta_{C}$. The vertical dashed lines labels $\beta_{C}$ we pinpoint for $t^{\prime}=0.0$ through finite-size analysis (inset). Here, $t=1$, $U=0.4$, $J=0.0$, $t^{\prime}=0.0,0.5$, and $L=70$. We provide consistent MC and exact results for $t^{\prime}=0$, whereas only MC results are available at $t^{\prime}=0.5$.
Long-range order and energetics— We numerically evaluate physical properties such as the expectation value $\langle X\rangle=\langle|\sum_{i}X_{i}|\rangle/L$ and the specific heat $C_{V}$ over various system sizes and observe clear signatures of ordered phases at sufficiently low temperatures ($\beta>\beta_{C}$); see examples in Fig. 2. Here, $\langle X\rangle$ serves as an Ising-type order parameter and its long-range correlation emerges in the ordered phases [78]. The finite-size scaling of the specific heat $C_{V}$ accurately locates the transitions. These transitions exhibit first-order characteristics, as evident from the abrupt bifurcation of the lowest-free-energy configuration from zero to two opposite values in $X$ [78], accompanied by spontaneous breaking of the $D$ symmetry through the selection of either of the two states with positive or negative $X$. We summarize the results in the phase diagrams in Fig. 3, where the ordered phases hold finite regions leaning towards larger $U$ and lower $T$, indicating their sufficient robustness towards variations of parameters. Beyond the primary transition, the system exhibits a sequence of phase transitions with distinctive signatures in both $\langle X\rangle$ and $C_{V}$, whose physical mechanism will be explained afterward. On the contrary, the entire parameter spaces consist of only a disordered phase under open boundary conditions (OBCs) or without non-Hermitian construction, e.g., $U=0$ or $U=0.4i$ [78].
@@
Let us consider the $t^{\prime}=0$ case in Eq. 1 as an example: given a specific $X$ configuration, the ground-state energy $E(X)$ of $\hat{H}(X)$ depends only on $n_{\pm}$; the minimum $E(X)$ corresponds to the uniform configuration $X_{0}$ with all $X_{i}=1$ ($X_{i}=-1$). Importantly, unlike the Ising interaction that contributes a constant energy cost of $2J$ for each domain wall, $E(X)$, through non-Hermitian construction, is no longer locally dependent on the (number of) domain walls in $X$, but on the separations between them as well. For instance, $E(X)$ increases linearly with two domain walls $r=\alpha L$ apart, while $E(X)$ saturates to a constant with nearby domain walls in the thermodynamic limit, as we illustrate in Fig. 4. The trend remains at finite temperature. Such non-local energetic dependence applies to general scenarios beyond similarity transformations, e.g., $t^{\prime}\neq 0$, but are absent in similar yet Hermitian constructions or under OBCs [78].
Figure 4: The ground-state energy $E(X)$ increases linearly with two domain walls $r=\alpha L$ apart and saturates to a constant with nearby domain walls in the thermodynamic limit. Note the differences between the vertical scales. Here, $t^{\prime}=0,0.5$, $U=0.4$, and $X_{0}$ is uniformly ordered without domain walls. The insets display schematic diagrams of the corresponding $X$ configurations featuring two domain walls, whose energy is attributed to both the fermions and the Ising interaction (a constant $2J$ energy cost) upon $X$.
@@
Topological aspects of fermions— Here, we examine the SSB and ordered phases from a fermionic perspective. The semiclassical theory has suggested that the complex energy spectrum in non-Hermitian systems may induce a residual imaginary velocity [35, 36]. In Fig. 5, we numerically compute the statistical distribution of such imaginary velocities, which exhibits a behavior analogous to the Ising-model magnetization across its SSB: (i) a peak at zero in the disordered phase and (ii) spontaneous bifurcation into conjugate imaginary sectors (±$\pm\operatorname{Im}v$) in the ordered phases. Moreover, long-range correlations of local $v$ emerge in the ordered phases. While the global Hermiticity requires $\langle v\rangle=0$, $\langle|v|\rangle$ emerges as a natural order parameter.
@@
Figure 5: The phases and phase transitions exhibit a topological nature from the fermion perspective. (a): The probability distribution of $v$ shifts from peaking at 0 to finite-imaginary-value concentrations. (b): Long-range correlation of local $v$ emerges in the ordered phases. (c): The fermion path integral’s winding number $w$ transitions from the conventional $|w|=0$ plateau (disordered phase) to quantized plateaus $|w|=2,4,\cdots$ (ordered phase). The vertical dashed line labels the corresponding transitions in Fig. 2; temperatures in (a-b) align with colored arrows in (c). Here, $U=0.4$, $J=0.0$, and $t^{\prime}=0.0$. Inset in (c): (Upper right) an illustration of nontrivial winding number $w=2$ in (1+1)D path integral under PBCs, and (Lower left) an illustration of the imaginary velocity as from the complex energy spectrum, detailed in [78, 36].
In quantum Monte Carlo simulations depicting imaginary time evolution within the (1+1)D path integral framework, the worldline winding number $w$ counts the number of times that the particle trajectories (worldlines) cross the periodic boundaries during an imaginary time $\beta=1/k_{B}T$. Conventionally, in Hermitian systems, the most relevant worldlines reside predominantly in the $w=0$ sector [81]; however, this no longer holds for non-Hermitian systems under PBCs: the dominant worldlines winding number locates in $w\neq 0$ sector and is related to the imaginary velocity through $w=\beta\operatorname{Im}(\langle v\rangle)$  [^3] This equation is taken from a previous result [36], with the parameter $L$ absorbed into our definition of $\langle v\rangle$; see the Supplemental Material [78] for details. , which reflects that the quantum dynamics of imaginary velocity induce unidirectional collective displacements in imaginary time [78, 35, 36]. Here, we note that the SSB is also signaled and accompanied by the shift of dominance from the $w=0$ sector to the $w\neq 0$ sectors: while the disordered phase corresponds to a $|w|=0$ plateau, multiple order phases agree with respective nontrivial integer plateaus of winding numbers $|w|=2,4,\cdots$, separated by abrupt changes at critical temperatures consistent with the phase transitions from $\langle X\rangle$ and $C_{V}$ in Fig. 2. The even quantization in $|w|$ stems from the independent and simultaneous left and right Fermi-point contributions, and can be modified by de-synchronizing the Fermi-point contributions - for example, adding time-reversal symmetry-breaking terms to Eq. 1 yields $\Delta|w|=\pm 1$ plateaus, see the Supplemental Material for details [78].
As the $D$ symmetry is intimately related to Hermiticity, the system exhibits non-Hermitian characteristics after SSB, including finite imaginary velocity and non-zero winding numbers. It also leaves nontrivial marks on the fermion spectral functions: while singularities mainly concentrate on the real axis on the complex energy plane in the disordered phase, they materialize in imaginary regions on the complex energy plane in the order phases [78]. While spontaneous “Hermiticity breaking” might appear unphysical, it bears a profound analogy to “gauge-symmetry breaking” in superconductors - the “breaking” is only explicit at a mean-field level, as exemplified by a finite $\langle X\rangle$ in Eq. 5 [78].
@@
Acknowledgment: We acknowledge insightful discussions with Yongxu Fu and Shi-Xin Hu, and generous support from the National Natural Science Foundation of China (Grants No.92270102 & No.12174008) and the National Key R&D Program of China (Grant No.2022YFA1403700).